gelphielvr:

(pre geometry?) 3x+2 2x+1 5x-1 ------- + ------- = ------- 3 2 6

gelphielvr:

I got x = 2 but I was told it was wrong but I don't know where i did my mistake

Spectrum:

huh? do you have an image

gelphielvr:

@spectrum wrote:
huh? do you have an image
it was on paper so no - the ---- is supposed to be a fraction cause I didn't know how else to illustrate it

Spectrum:

so does it looks somewhat like this? (Give me a minute to type it)

Spectrum:

\[\frac{3x+2}{3}+\frac{2x+1}{2}=\frac{5x-1}{6}\]

gelphielvr:

yes exactly like that sorry I didn't know how to make it look like that.

Spectrum:

\[6*(\frac{3x+2}{3}+\frac{2x+1}{2})=6*(\frac{5x-1}{6})\]

Spectrum:

we do this bc 6 is the least common multiple of the denominators (6), since there's a 6, 3, ad 2.

Spectrum:

for the right side, we are not changing it, since it was ____ over 6, multiplying it by 6 just canceled out the fraction.

Spectrum:

\[\frac{6}{3}*(3x+2)\]

Spectrum:

@spectrum wrote:
\[6*(\frac{3x+2}{3}+\frac{2x+1}{2})=6*(\frac{5x-1}{6})\]
we basically just moved the 3 out from the \[\frac{3x+2}{3}\] and put it below the 6

Spectrum:

@spectrum wrote:
\[\frac{6}{3}*(3x+2)\]
which becomes \[2*(3x+2)\]

Spectrum:

and then doing the other side we get \[3*(3x+1)\]

Spectrum:

then \[2*(3x+2)+3*(3x+1)=5x-1\]

Spectrum:

alright, now we do the distributive property.

Spectrum:

\[6x+4+6x+3=5x-1\]

Spectrum:

combine like terms to get: \[12x+7=5x-1\]

Spectrum:

move the 5x over, you want the variables to be on the left side. \[12x-5x+7=5x-5x-1\]

Spectrum:

@spectrum wrote:
move the 5x over, you want the variables to be on the left side. \[12x-5x+7=5x-5x-1\]
\[7x+7=-1\]

Spectrum:

subtract 7 from both sides since you need a number on the right side, this gives you \[7x=-8\]

Spectrum:

and divide both sides by 7 so you get the x alone \[x=\frac{-8}{7}\]

Spectrum:

so that's your answer. \[x=\frac{-8}{7}\]